OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

-9.1° 1.8° 0.9° 11.8°

OpenStudy (anonymous):

\[u\cdot v=|u||v|\cos\theta\]